40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 30 13:02:51 2013, Masayuki,Manasa, Update, IOO, PMC and MC alignment 
    Reply  Thu Oct 3 01:24:31 2013, Jenne, Rana, Update, IOO, PMC  PMCR_1064822387.bmppmc80.png
       Reply  Sun Oct 6 22:52:09 2013, Jenne, Rana, Update, IOO, input beam to PMC aligned again pmcr.png
          Reply  Mon Oct 14 17:51:40 2013, Jenne, Update, IOO, input beam to PMC drifted again 
             Reply  Mon Oct 21 13:15:52 2013, Koji, Update, IOO, PMC aligned 
             Reply  Fri Oct 25 10:45:16 2013, Steve, Update, PSL, laser drift monitor set up laserDriftMon.jpgPMCT_120d.pngPMCT_1000d.png
                Reply  Fri Oct 25 19:56:58 2013, rana, Update, PSL, laser drift monitor set up PMC-IRSISSS.pngPA250052.JPGPA280044.JPG
                Reply  Fri Jan 10 15:33:02 2014, Steve, Update, PSL, laser drift monitor set up idea 2daysDrift.png
                   Reply  Mon Jan 13 11:08:48 2014, Steve, Update, PSL, 3 good days of IOO pointing 3gdPSLpointing.png
                   Reply  Tue Jan 14 10:12:12 2014, Steve, Update, PSL, laser drift monitor set up idea bad4thday.png
Message ID: 9292     Entry time: Fri Oct 25 19:56:58 2013     In reply to: 9291
Author: rana 
Type: Update 
Category: PSL 
Subject: laser drift monitor set up 

 

 I went to re-align the beam into the PMC just now. I also tapped all the components between the laser and the PMC; nothing seems suspicious or loose.

The only problem was that someone (probably Steve or Valera) had closed down the iris just downstream of the AOM to ~1-2 mm diameter. This is much too tight! Don't leave irises closed down after aligning. An iris is not to b used as a beam dump. Getting it within a factor of 5-10 of the beam size will certainly make extra noise from clipping/scattering. After opening the iris, the reflected beam onto the PMC REFL camera is notably changed.

Not sure if this will have any effect on our worsening transmission drift, but let's see over the weekend.

I took pictures of this clipping as well as the beam position on Steve's new Retro Position Sensor, but I can't find the cable for the Olympus 570UZ. Steve, please buy a couple more USB data cables of this particular kind so that we don't have to hunt so much if one of the cryo (?) people borrows a cable.

Attachment shows PMC power levels before and after alignment. After alignment, you can see spikes from where I was tapping the mounts in the beamline. We ought to replace the U-100 mount ahead of the AOM with a Polanski

EDIT: Cryo team returns cable - receives punishments. Picture added.

Attachment 1: PMC-IRSISSS.png  21 kB  | Show | Hide all | Show all
Attachment 2: PA250052.JPG  716 kB  Uploaded Mon Oct 28 22:04:53 2013  | Show | Hide all | Show all
Attachment 3: PA280044.JPG  303 kB  Uploaded Mon Oct 28 22:17:04 2013  | Hide | Hide all | Show all
PA280044.JPG
ELOG V3.1.3-